the length of a rectangle field is 200 m.if its perimeter is 700 m.find its breadth
Answer by Cromlix(4381)

Perimeter = 2 x length + 2 x breadth
Therefore 2 x 200m + 2 x breadth = 700 m
400 m + 2b = 700 m
2b = 700m - 400m
2b = 300 m
Therefore breadth = 300/2 = 150 m
Hope this helps :-)